# Is UEFA really committed to sporting merit? A misaligned rule in the post-2024 Champions League

### v3

@@ Abstract
UEFA declares that it is fully committed to respecting the fundamental values of sport. However, the qualification of the post-2024 UEFA Champions League is not always based on sporting merit. Consequently, a match with misaligned incentives has been barely avoided in the 2023/24 German Bundesliga. Two straightforward solutions are recommended to safeguard the integrity of the competition. Fairness can be easily guaranteed by reversing the order of two articles in the Champions League regulation. Since UEFA repeatedly commits the same mistake that is widely known among researchers, governing bodies in sports are called to work more closely together with the scientific community.
@@ 1 Introduction
Ensuring the fairness and integrity of competitions is a key responsibility of sports policy (Sugden et al.,, 2023). Naturally, the Union of European Football Associations (UEFA) is also committed to these values, at least in declarations and media releases. For example, UEFA president Aleksander Čeferin has commented on the novel format of the UEFA Champions League with the following words (UEFA, 2024a, ):
@@
However, is the UEFA Champions League qualification based on sporting merit under any circumstances? Is there any issue in its rules that compromises the fairness of this tournament and can potentially jeopardise competition?
The current note uncovers that, while UEFA may be committed to sporting merit in principles, their rules do not avoid an unfair game in a domestic league. In particular, although the recent fundamental reform of the UEFA Champions League has restored the integrity of the seeding system, the way in which the novel European performance spots are assigned allows for a situation when a team is strictly better off by losing. These misaligned incentives almost occurred in the last round of the 2023/24 German Bundesliga in the match Eintracht Frankfurt vs. RB Leipzig.
This problem is especially worrying because “the objective for a sport organisation should be to eliminate any game that raises potential unfairness ex-ante(Scelles et al.,, 2024, p. 535). Furthermore, its solution has been known since March 2013 when the first version of Dagaev and Sonin, (2018) become available.
@@ 2 Related literature
This theoretical finding might seem not especially interesting for sports policy. However, the misaligned allocation rule has been responsible for the incentive incompatibility of
@@ 4 Discussion and policy implications
A recent literature survey on fraud in sport (Vanwersch et al.,, 2022) concludes that future research is needed to accumulate more knowledge on the prevention of fraud. Our paper has examined the new format of the UEFA Champions League from this perspective. It should be acknowledged that the rules have improved from the 2024/25 season. Now the clubs are seeded based on their individual club coefficients except for the guaranteed place of the titleholder in the first pot (UEFA, 2024b, , Article 16.01). Therefore, the regime used until the 2014/15 season has returned, and the seeding system satisfies incentive compatibility.
@@
Naturally, the set of matches to be played cannot be changed in a domestic championship without changing the tournament format, but the order of the matches is a decision variable. The scenario outlined in Section 3 has become threatening especially since the match Eintracht Frankfurt vs. RB Leipzig is played in the last round. The set of top teams can be relatively accurately estimated at the beginning of the season, hence, the last rounds can be scheduled to avoid any matches between the strongest teams. Such a schedule may lead to a tanking opportunity at the end of the season with a (much) lower probability. However, even if Eintracht Frankfurt had already played against RB Leipzig, it would be unfair if it turns out ex post that a team has missed qualifying for the UEFA Champions League due to winning a match months before. Therefore, it is the responsibility of UEFA to act if it is really committed to its announced principles.
@@ 5 Conclusions
  • All proposed rule changes should be posted in a public forum, where interested academics could comment, perhaps after running simulations.
Hopefully, UEFA and all other sports governing bodies will adopt one of these options in the future. This would be especially important because, as it has been highlighted in Section 2, UEFA repeatedly commits the same mistake in allocating vacant slots—while the fair solution has been provided by Dagaev and Sonin, (2018), and the first version of this paper is available since March 2013.

### v4

# UEFA Champions League entry is incentive incompatible again from the 2024/25 season

@@ Abstract
UEFA declares that it is fully committed to respecting the fundamental values of sport. However, the qualification for the post-2024 UEFA Champions League is not always based on sporting merit. Consequently, a match with misaligned incentives has been barely avoided in the 2023/24 German Bundesliga. Two straightforward solutions are recommended to safeguard the integrity of the competition. Fairness can be easily guaranteed by reversing the order of two articles in the Champions League regulation. Since UEFA repeatedly commits the same mistake that is widely known among researchers, governing bodies in sports are called to work more closely together with the scientific community.
@@ 1 Introduction
Ensuring the fairness and integrity of competitions is a key responsibility of tournament organisers (Kendall and Lenten,, 2017; Devriesere et al.,, 2024). Naturally, the Union of European Football Associations (UEFA) is also committed to these values, at least in declarations and media releases. For example, UEFA president Aleksander Čeferin has commented on the novel format of the UEFA Champions League in 2024 with the following words (UEFA, 2024a, ):
@@
However, is the UEFA Champions League qualification based on sporting merit under any circumstances? Is there any issue in its rules that threatens the fairness of this tournament and can potentially jeopardise competition?
The current note uncovers that, while UEFA may be committed to sporting merit in principles, their regulations do not avoid an unfair game in a domestic league. In particular, although the recent fundamental reform of the UEFA Champions League has restored the integrity of the seeding system, the way in which the novel European performance spots are assigned allows for a situation when a team is strictly better off by losing. These misaligned incentives almost occurred in the last round of the 2023/24 German Bundesliga in the match Eintracht Frankfurt vs. RB Leipzig.
The problem is especially worrying because “the objective for a sport organisation should be to eliminate any game that raises potential unfairness ex-ante(Scelles et al.,, 2024, p. 535). Furthermore, its solution has been known since March 2013 when the first version of Dagaev and Sonin, (2018) became available.
@@ 2 Related literature
This theoretical finding might not seem especially interesting for sports policy. However, the misaligned allocation rule has been responsible for the incentive incompatibility of
@@ 4 Discussion and policy implications
A recent literature survey on fraud in sport (Vanwersch et al.,, 2022) concludes that future research is needed to accumulate more knowledge on the prevention of fraud. Our paper has examined the new format of the UEFA Champions League from this perspective. The rules have clearly improved from the 2024/25 season. Now the clubs are seeded based on their individual club coefficients except for the guaranteed place of the titleholder in the first pot (UEFA, 2024b, , Article 16.01). Therefore, the regime used until the 2014/15 season has returned, and the seeding system satisfies incentive compatibility.
@@
Naturally, the set of matches to be played cannot be changed in a domestic championship without changing the tournament format, but the order of the matches is a decision variable. The scenario outlined in Section 3 has become threatening especially since the match Eintracht Frankfurt vs. RB Leipzig has been played in the last round. The set of top teams can be relatively accurately estimated at the beginning of the season, hence, the last rounds can be scheduled to avoid any matches between the strongest teams. Such a schedule may lead to a tanking opportunity at the end of the season with a (much) lower probability. However, even if Eintracht Frankfurt had already played against RB Leipzig, it would be unfair if it turns out ex post that a team has missed qualifying for the UEFA Champions League due to winning a match months before. Therefore, it is the responsibility of UEFA to act if it is really committed to its announced principles.
@@ 5 Conclusions
  • All proposed rule changes need to be posted in a public forum, where interested academics could comment, perhaps after running simulations.
Hopefully, UEFA and other sports governing bodies will adopt one of these options in the future. This would be especially important because, as it has been highlighted in Section 2, UEFA repeatedly commits the same mistake in allocating vacant slots—while the fair solution has been provided by Dagaev and Sonin, (2018), and the first version of this paper is available since March 2013.